OpenStudy (anonymous):

how do you find a side length of a 45 45 90 triangle

OpenStudy (anonymous):



OpenStudy (unklerhaukus):

pythagorus says that \[a^2+b^2=c^2\] where \(a\) and \(b\) are both \(8\) and \(c\) is \(x\)

OpenStudy (anonymous):

so then a^2+b^2 is 128 i square that to get my answer or is it just 128

OpenStudy (unklerhaukus):

128 = c^2 take the square root the get c √ 128 =c

OpenStudy (anonymous):

thank you

OpenStudy (unklerhaukus):

what do you get as you final answer?

OpenStudy (anonymous):

11.31

OpenStudy (triciaal):

 for a 45-45-90 triangle the legs are equal let =x use pythagorean theorem as given by !UncleRhaukus above